# Improving Multi-candidate Speculative Decoding

### v2

@@ 3.1 Target Model Initialized Multi-Candidate Generation
Existing multi-candidate speculative decoding methods rely on a draft model to generate the entire multi-candidate token tree, and after verifying the draft generated tokens, only sample one token from the target model or normalized target and draft model’s output distribution. Due to the difference between the output distribution of the draft and target model, there is no specific criterion to determine which token sampled from the target model will yield the longest accepted token sequence for future draft model generation. Therefore, we hypothesize that sampling multiple tokens instead of one token from the target model’s distribution to initialize a multi-candidate sequence for future draft model generation can increase the acceptance rate. However, When the target model initializes more than one token for multi-candidate generation, selecting an initialized token based on the longest accepted draft token sequence creates a dependency. This means that the acceptance probabilities of the sequential draft tokens influence the probability of accepting the target-sampled token. Consequently, the output distribution no longer aligns with the target model’s. Specifically, by selecting the target token $t$ that yields the longest accepted draft sequence, the probability of accepting the target sampled token $t$ becomes:
$P_{\text{output}}(t)=P_{\text{target}}(t)\times P_{\text{acceptance}}(t+1,t+2,\dots\mid t)$
Where $P_{\text{acceptance}}(t+1,t+2,\dots\mid t)$ is the probability that the subsequent draft tokens are accepted given the target token $t$. This alters the output distribution, causing it to diverge from $P_{\text{target}}(t)$. Thus, we evaluate the quality of our method’s output using MT-Bench. The empirical results indicate that the quality loss depends on the number of target model initialized tokens and the draft model. For more details, please refer to the experimental section.
@@ 3.3 Decision Model
We design two types of decision models to dynamically determine if the early stop is necessary for the draft generation process. The first type of decision model is a three-layer MLP, taking the hidden states from draft model as input and the $max(1,\frac{p(x|I)}{q(x|I)})$ ($p(x|I)$ is the probability of target model predict token $x$ with given input $I$ and $q(x|I)$ is the probability of draft model predict token $x$ with given input $I$) as the training label. The decision process can be represented as:
@@ 4.2 Overall Results
Target Initialized Width TinyLlama-1.1B Llama-68M

### v3

@@ 3.1 Target Model Initialized Multi-Candidate Generation
Existing multi-candidate speculative decoding methods rely on a draft model to generate the entire multi-candidate token tree, and after verifying the draft generated tokens, only sample one token from the target model or normalized target and draft model’s output distribution. Due to the difference between the output distribution of the draft and target model, there is no specific criterion to determine which token sampled from the target model will yield the longest accepted token sequence for future draft model generation. Therefore, we hypothesize that sampling multiple tokens instead of one token from the target model’s distribution to initialize a multi-candidate sequence for future draft model generation can increase the acceptance rate. However, When the target model initializes more than one token for multi-candidate generation, selecting an initialized token based on the longest accepted draft token sequence creates a dependency. This means that the acceptance probabilities of the sequential draft tokens influence the probability of accepting the target-sampled token. Consequently, the output distribution no longer aligns with the target model’s. Specifically, by selecting the target token $t_{i}$ that yields the longest accepted draft sequence, the probability of accepting the target sampled token $t_{i}$ becomes:
$P_{\text{output}}(t_{i})=P_{\text{target}}(t_{i})\times P_{\text{acceptance}}(t_{i+1},t_{i+2},\dots\mid t_{i})$
Where $P_{\text{acceptance}}(t_{i+1},t_{i+2},\dots\mid t_{i})$ is the probability that the subsequent draft tokens are accepted given the target token $t_{i}$. This alters the output distribution, causing it to diverge from $P_{\text{target}}(t_{i})$. Thus, we evaluate the quality of our method’s output using MT-Bench. The empirical results indicate that the quality loss depends on the number of target-driven initialization tokens and the draft model. For more details, please refer to the experimental section.
@@ 3.3 Decision Model
We design two types of decision models to dynamically determine whether an early stop is necessary during the draft generation process. The first type of decision model is a three-layer MLP, taking the hidden states from draft model as input and the $max(1,\frac{p(x|I)}{q(x|I)})$ ($p(x|I)$ is the probability of target model predict token $x$ with given input $I$ and $q(x|I)$ is the probability of draft model predict token $x$ with given input $I$) as the training label. The decision process can be represented as:
@@ 4.2 Overall Results
Target Initialized Width Llama-68M TinyLlama-1.1B